Sep. 3 Purchased merchandise inventory on account from Sheldon Wholesalers, $ 3,500. Terms 1 / 15, n / E O M, FOB shipping point.

4 Paid freight bill of $ 75 on September 3 purchase.

4 Purchase merchandise inventory for cash of $ 1,900.

6 Returned $ 400 of inventory from September 3 purchase.

8 Sold merchandise inventory to Harvey Company, $ 6,300, on account. Terms 1 / 15, n / 35. Cost of goods, $ 2,835

9 Purchased merchandise inventory on account from Tripp Wholesalers, $ 12,000. Terms 1 / 10, n / 30, FOB destination.

10 Made payment to Sheldon Wholesalers for goods purchased on September 3 , less return and discount.

12 Received payment from Harvey Company, less discount.

13 After negotiations, received a $ 100 allowance from Tripp Wholesalers.

15 Sold merchandise inventory to Johnson Company, $ 3,500, on account. Terms n / E O M. Cost of goods, $ 1,400.

22 Made payment, less allowance, to Tripp Wholesalers for goods purchased on September 9 .

23 Johnson Company returned $ 800 of the merchandise sold on September 15. Cost of goods, $ 320.

25 Sold merchandise inventory to Smecker for $ 1,500 on account that cost $ 555 . Terms of 1 / 10, n / 30 was offered, FOB shipping point. As a courtesy to Smecker, $ 90 of freight was added to the invoice for which cash was paid by Cardinal.

29 Received payment from Smecker, less discount.

30 Received payment from Johnson Company, less return.


Sep. 3 4 4 6 8 Purchased merchandise inventory on account from Sheldon Wholesalers, $3,500. Terms 1/15, n/EOM, FOB shipping pTL TC U UISCOUTIC 12 13 15 Received payment from Harvey Company, less discount. After negotiations, received a $100 allowanceSep 3: Purchased merchandise inventory on account from Sheldon Wholesalers, $3,500. Terms 1/15, n/EOM, FOB shipping point. DaSep. 4: Paid freight bill of $75 on September 3 purchase. Date Accounts Debit Credit Sep. 4 Merchandise Inventory CashSep. 4: Purchase merchandise inventory for cash of $1,900. Date Accounts Debit Credit Sep. 4 Merchandise Inventory CashSep. 6: Returned $400 of inventory from September 3 purchase. Date Accounts Debit Credit Sep. 6 Accounts Payable—Sheldon Whol
Answer #1
Journal entries
Date Accounts Name Debit Credit Calculation
3-Sep Merchandise Inventory         3,500
Accounts Payable - Sheldon Wholesalers         3,500
4-Sep Merchandise Inventory               75
Cash               75
4-Sep Merchandise Inventory         1,900
Cash         1,900
6-Sep Accounts Payable - Sheldon Wholesalers            400
Merchandise Inventory            400
8-Sep Accounts Receivable - Harvey Company         6,300
Sales Revenue         6,300
Cost of goods sold         2,835
Merchandise Inventory         2,835
9-Sep Merchandise Inventory      12,000
Accounts Payable - Tripp Wholesalers      12,000
10-Sep Accounts Payable - Sheldon Wholesalers         3,100 =3,500-400
Cash         3,069
Merchandise Inventory               31 =3,100*1%
12-Sep Cash         6,237
Sales Discounts               63 =6,300*1%
Accounts Receivable - Harvey Company         6,300
13-Sep Accounts Payable - Tripp Wholesalers            100
Merchandise Inventory            100
15-Sep Accounts Receivable - Johnson Company         3,500
Sales Revenue         3,500
Cost of goods sold         1,400
Merchandise Inventory         1,400
22-Sep Accounts Payable - Tripp Wholesalers      11,900 =12,000-100
Cash      11,900
23-Sep Sales Returns & Allowances            800
Accounts Receivable - Johnson Company            800
Merchandise Inventory            320
Cost of goods sold            320
25-Sep Accounts Receivable - Smecker         1,590 =1,500+90
Sales Revenue         1,500
Cash               90
Cost of goods sold            555
Merchandise Inventory            555
29-Sep Cash         1,575
Sales Discounts               15 =1,500*1%
Accounts Receivable - Smecker         1,590
30-Sep Cash         2,700 =3,500-800
Accounts Receivable - Johnson Company         2,700